# Advancing Language Multi-Agent Learning with Credit Re-Assignment for Interactive Environment Generalization

### v2

@@ 4.1 Experimental Settings
Base Model GPT-3.5-Turbo N/A 1 Text 17.4 10.7 -
Base Model GPT-4 N/A 1 Text 30.9 37.8 -
Base Model Claude2 N/A 1 Text - 10.5 -
@@
Base Model LLaMA2 7B 1 Text - 2.9 -
Base Model LLaMA2 70B 1 Text - 10.6 -
Base Model Qwen2 7B 1 Text 07.4 8.5 -
SFT Qwen VL 9.6B 1 Text & Image 10.1 - -
SFT SeeClick 9.6B 1 Text & Image 20.9 - -
@@
Base UIAgent Qwen2 7B 1 Text 11.9 12.8 04.3
Base UIAgent Qwen2 7B 4 Text 13.2 14.0 05.5
CollabUIAgentsmobile Qwen2 7B 4 Text 16.2 17.7 09.2
@@
Table 2: Experimental results on web browsing environments. Average step success rates (SSR) in Mind2Web and AutoWebBench are reported. “SFT” denotes the base model is supervised fine-tuned in Mind2Web. $\boldsymbol{\Delta}$Generalization indicates the gap between the base model and agent learning methods based on the model in AutoWebBench.
4.2 Main Results
Effectiveness in Mobile Environments  In this section, we explore the effectiveness of our proposed method in trained and unseen mobile operation environments. Experimental results in mobile environments are shown in Table 1. The performance of proprietary models face instability with input format, e.g., text-only or multi-modal input. With appropriate agentic fine-tuning, open-source models like Qwen2 7B could gain significant performance improvement, and even higher within multi-agent systems (“Base UIAgent”). CollabUIAgentsmobile achieves the best results among systems based on open-source LLMs. Remarkably, it outperforms Gemini 1.5 Pro in both environments and achieves performance comparable to or better than GPT-4, due to its instability. These outcomes demonstrate the effectiveness of our framework and provide evidence for the validity of the CR strategy. On the other hand, InfiGUIAgent uses a large amount of training data but falls behind other methods; DigiRL shows good performance in the training environment, but falls short in generalization compared to CollabUIAgentsmobile, which improves greatly on unseen tasks.
@@ 4.2 Main Results
Base UIAgent 18.9 48.4
@@
Base UIAgent 21.4 53.2
@@
4.3 Ablation Study
(1) Further training of the Base UIAgent with trajectory data using either rejective SFT (“CollabUIAgentsmobile w/ PO $\rightarrow$ RFT”) or DPO on whole trajectories (“CollabUIAgentsmobile w/o CR”) improves performance, with DPO showing superior results. The primary distinction between these methods is that SFT can only learn from correct actions, while DPO can learn from both correct and incorrect actions. But they all faces reward scarcity problems, resulting in modest improvement. (2) CollabUIAgentsmobile introduces credit re-assignment, providing more granular feedback that facilitates exploration of the large action space at each step. The synthesized preference data also helps generalization through preference optimization. This boosts both performance and generalizability, yielding the best overall results.
(3) Combining multiple agents based on a vanilla base model using random edges leads to modest improvements, and the similar trend exists for DigiRL in Table 1, underscoring the importance of proper agentic fine-tuning, and MARL with online multi-agent trajectories. (4) A comparison between systems with and without edge updates (“CollabUIAgentsmobile” vs. “w/o edge update”) demonstrates that the edge update trick contributes to further accommodating language agents in complex multi-agent systems with communications.
@@ 5 Related Work
Agents on Interactive Environments  Before the advent of LLMs, agents relied on traditional RL to perform interactions such as clicking and typing (Liu et al., 2018b; Humphreys et al., 2022). However, recent advancements have shifted towards leveraging foundation models with in-context learning or fine-tuning across various interfaces, including mobile (Wang et al., 2023; Hong et al., 2024b), web (Lai et al., 2024; Deng et al., 2024b), and computer using environments (Xu et al., 2024; Wu et al., 2024b). Recently, there are emerging methods designing process rewards (He et al., 2024a; Pan et al., 2024; He et al., 2025; Huang et al., 2025) or language RL (Bai et al., 2024) for better performing single agents.
@@ 6 Conclusion
Appendix A Case Studies on Mobile Operation Environments
@@
Appendix B Agentic Fine-Tuning in CollabUIAgents
Figure 4: An example of task execution steps. “action_type” represents the action taken, and “index” represents the index of the UI element. The positions of the relevant elements on the UI interface are marked in red.
B.1 Methodology
@@
(1) Basic Environmental Knowledge: This data segment includes identifying UI elements and understanding their properties. We categorize basic knowledge into two types: UI Understanding (coarse-grained): This refers to a broad understanding of the layout and information contained in the UI, such as identifying the purpose of the interface. UI Element Recognition (fine-grained): Since UI typically contains a large number of densely packed interface, the agent needs to be able to distinguish between different types of elements, such as buttons, input fields, and drop-down menus, and understand the associated actions. We develop a series of queries accordingly in Appendix C.4.1, and randomly select UI elements and the layout to assemble queries for the UI Agent.
@@
(3) Process Preference Knowledge: Real-world interactive tasks is quite difficult, and even the most advanced large language model, GPT-4, shows a low task completion rate (30%) in the mobile environment AndroidWorld (Rawles et al., 2025). Training a model solely on scarce successful trajectories still inevitably results in errors. Therefore, as illustrated below Figure 3, we introduce the adversarial agent against the UI agent, and the critic agent to score all actions, obtaining process preference data with step-level rewards. By learning from process preference data, the agent can better distinguish between correct and incorrect actions during the process, ultimately improving task completion rates. The distribution of the collected data can be found in Appendix C.4.2.
@@
B.2 Ablation Study of Agentic Fine-tuning
@@
Appendix C Experiment Details
C.1 Action Space in the Environments
Table 5 and 6 show the action spaces of agents in mobile and web environments, respectively.
@@ C.1 Action Space in the Environments
C.2 Detailed Results in Web Browsing Environments
@@
System #Params #Agents Input Cross-Task Cross-Website Cross-Domain Avg.
@@
GPT-3.5-Turbo N/A 1 Text 17.4 16.2 18.6 17.4
GPT-4 N/A 1 Text 36.2 30.1 26.4 30.9
@@
Qwen-VL* 9.6B 1 Text & Image 12.6 10.1 08.0 10.2
SeeClick* 9.6B 1 Text & Image 23.7 18.8 20.2 20.9
Qwen2 7B 1 Text 08.6 06.3 07.5 07.4
@@
Base UIAgent 7B 1 Text 13.4 10.6 11.8 11.9
Base UIAgent 7B 4 Text 15.7 11.2 12.9 13.2
CollabUIAgentsmobile 7B 4 Text 19.2 13.8 15.5 16.2
@@
System #Params #Agents English Chinese Avg.
@@
LLaMA2 7B 1 03.3 02.5 - - 02.9
LLaMA2 70B 1 08.3 08.9 - - 10.6
Qwen2 7B 1 08.6 09.4 08.1 07.8 08.5
@@
Base UIAgent 7B 1 12.0 13.3 12.7 13.4 12.8
Base UIAgent 7B 4 13.7 14.5 15.0 13.9 14.0
CollabUIAgentsmobile 7B 4 18.6 17.7 19.1 15.6 17.7
@@
C.3 Method Implementation
C.3.1 Baselines
@@
(1) M3A (Rawles et al., 2023) is a prompt-based multimodal agent, which combines ReAct- (Yao et al., 2023) and Reflexion-style (Shinn et al., 2023) prompting to interpret user instructions and screen content, then update its decisions. (2) SeeAct (Zheng et al., 2024) is a prompt-based navigation agent originally designed for GPT-4V to perform actions with visual input and textual choices. (3) InfiGUIAgent (Liu et al., 2025) is used with the temperature set to 0.1, and other settings remain default. (4) DigiRL (Bai et al., 2024) is reproduced on the same dataset as CollabUIAgents, based on Qwen-2.5-VL-3B-Instruct (Wang et al., 2024d). Due to limited computational resource and the high costs of RL training with multiple value estimators, a larger scaled base model could not be applied. (5) SeeClick (Cheng et al., 2024b) is a fine-tuned visual GUI agent that automates tasks relying on screenshots and employs GUI grounding. For those baselines based on closed-source LLMs, we conduct experiments using prompts and settings consistent with Rawles et al. (2025), with model data up to August 2024.
C.3.2 CollabUIAgents Framework
@@
In the agentic fine-tuning process for data collection, we employ GPT-4o-2024-08-06 (OpenAI, 2024) as the UI agent. We set the temperature of different agents to 0.1 to ensure the accuracy of their responses. For the backbone model in the policy multi-agent system, we utilize Qwen2 7B (Yang et al., 2024). Detailed process is depicted in Appendix C.4. In this phase, the model undergoes both supervised learning and off-line preference learning with DPO. Details of the experimental data statistics can be found in Appendix C.4.2. During the supervised fine-tuning phase, the model’s context length is set to 8,192, with a learning rate of 1e-4 and training conducted over 3 epochs. During the preference optimization process, we adapt LoRA fine-tuning (Hu et al., 2022), and the model’s context length is capped at 8,500, with a learning rate of 5e-6. We use bf16 precision, and the training is also carried out for 3 epochs.
@@
In the multi-agent reinforcement learning process, the critic agent is played by GPT-4o-2024-08-06, with its temperature also set to 0.1, providing reward feedback to the rolled out actions. In the MARL phase, the rolling out temperatures of different UI agents are set to 0.1, 0.3, 0.5, and 0.8, respectively. This variation encourages diverse responses and stimulates different behaviors. We use iterative-DPO (Xiong et al., 2024) as the online reinforcement algorithm, and the multi-agent system is updated for each epoch. We train the multi-agent system within the AndroidWorld environment, where the environment seed is randomly generated. Continue MARL on Mind2Web uses instruction datasets autonomously collected with the pipeline in Appendix B. During testing, we maintain consistency with previous work (Rawles et al., 2025) by setting the seed to 30. All our experiments are conducted on 8 A100 80GB GPUs. The model’s context length is capped at 8,500, with a learning rate of 5e-6. We use bf16 precision, and the training is carried out for 3 epochs.
C.4 Data Collection for Agentic Fine-Tuning
@@ 4. What action is associated with UI element X?
C.4.2 Statistics of the Collected Data
@@ C.4.3 Prompts
C.5 Prompts for Different Agents in the CR
@@
C.6 Prompts for CollabUIAgents Framework

### v3

@@ 4.2 Main Results
GPT-3.5-Turbo N/A Text 17.4 10.7 -
GPT-4 N/A Text 30.9 37.8 -
Claude2 N/A Text - 10.5 -
@@
Qwen2 7B/1A Text 07.4 8.5 -
LLaMA2 7B/1A Text - 2.9 -
LLaMA2 70B/1A Text - 10.6 -
SFT (Qwen VL) 9.6B/1A Multimodal 10.1 - -
SeeClick (Qwen VL) 9.6B/1A Multimodal 20.9 - -
@@
UIAgent (Qwen2) 7B/1A Text 11.9 12.8 04.3
UIAgent (Qwen2) 7B/4A Text 13.2 14.0 05.5
CollabUIAgentsmobile (Qwen2) 7B/4A Text 16.2 17.7 09.2
@@
Table 2: Experimental results on web browsing environments. Average step success rates (SSR) in Mind2Web and AutoWebBench are reported. “SFT” denotes the base model is supervised fine-tuned in Mind2Web. $\boldsymbol{\Delta}$Generalization indicates the gap between the base model and agent learning methods based on the model in AutoWebBench. “7B/4A” denotes a four-agent system upon a 7B-parameter model.
Effectiveness in Mobile Environments  In this section, we explore the effectiveness of our proposed method in trained and unseen mobile operation environments. Experimental results in mobile environments are shown in Table 1. The performance of proprietary models face instability with input format, e.g., text-only or multi-modal input. With appropriate agentic fine-tuning, open-source models like Qwen2 7B could gain significant performance improvement, and even higher within multi-agent systems (“Base UIAgent”). CollabUIAgentsmobile achieves the best results among systems based on open-source LLMs. Remarkably, it outperforms Gemini 1.5 Pro in both environments and achieves performance comparable to or better than GPT-4, due to its instability. These outcomes demonstrate the effectiveness of our framework and provide evidence for the validity of the CR strategy. On the other hand, InfiGUIAgent uses a large amount of training data but falls behind other methods; DigiRL shows good performance in the training environment, but falls short in generalization compared to CollabUIAgentsmobile, which improves greatly on unseen tasks. In addition, increasing the number of agents leads to further improvement in the training environment but does not improve as much in unseen environments, showing a trade-off. This identifies the importance of maintaining the same number of agents in both the MARL and the deployment stages for the best effectiveness and generalization.
@@
4.3 Ablation Study
(1) Further training of the Base UIAgent with trajectory data using either rejective SFT (“CollabUIAgentsmobile w/ PO $\rightarrow$ RFT”) or DPO on whole trajectories (“CollabUIAgentsmobile w/o CR”) improves performance, with DPO showing superior results. The primary distinction between these methods is that SFT can only learn from correct actions, while DPO can learn from both correct and incorrect actions. (2) CollabUIAgentsmobile introduces credit re-assignment, providing more granular feedback that facilitates exploration of the large action space at each step. The synthesized preference data also helps generalization through preference optimization, compared to CollabUIAgentsmobile w/o CR which only uses the outcome reward $R_{o}$. This boosts both performance and generalizability, yielding the best overall results.
@@ 4.3 Ablation Study
UIAgent (LLaMA2) 15.1 43.7
UIAgent (Qwen2) 18.9 48.4
UIAgentself-critic (Qwen2) 10.7 19.5
@@
UIAgent (Qwen2) 21.4 53.2
UIAgentself-critic (Qwen2) 12.5 26.1
@@
(3) Combining multiple agents based on a vanilla base model using random edges leads to modest improvements (“Base UIAgent” ($n=4$)), and the similar trend exists for DigiRL in Table 1, underscoring the importance of proper agentic fine-tuning and MARL with online multi-agent trajectories. (4) A comparison between systems with and without edge updates (“CollabUIAgentsmobile” vs. “w/o edge update”) demonstrates that the edge update trick contributes to further accommodating language agents in complex multi-agent systems with communications.
@@ 5 Related Work
Agents on Interactive Environments  Before the advent of LLMs, agents relied on traditional RL to perform interactions such as clicking and typing (Liu et al., 2018b; Humphreys et al., 2022). However, recent advancements have shifted towards leveraging foundation models with in-context learning or fine-tuning across various interfaces, including mobile (Wang et al., 2023; Hong et al., 2024b), web (Lai et al., 2024; Deng et al., 2024b), and computer using environments (Xu et al., 2024a; Wu et al., 2024c). Recently, there are emerging methods designing process rewards (He et al., 2024a; Pan et al., 2024; Xu et al., 2024b; Wu et al., 2024b; He et al., 2025; Xu et al., 2025; Liu et al., 2025b), synthetic data (Yuan et al., 2024; Qin et al., 2025) and language RL (Bai et al., 2024) for better performing single agents.
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62276152), and funding from Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120.
Appendix A Comparisons with Non-Language Multi-Agent Learning Methods
To situate CollabUIAgents among recent studies that also employ LLMs for multi-agent credit assignment, we contrast it with the most closely related approaches of Lin et al. (2025); Qu et al. (2025). Table 4 summarizes the key differences.
Feature CollabUIAgents LCA (Lin et al., 2025) LaRe (Qu et al., 2025)
Primary focus Perf. & gen. of language agents across interactive environments Sparse-team credit assignment Episodic credit assignment (redundancy/ambiguity)
Core application Mobile/web dependent language tasks GridWorld, Pistonball MuJoCo, MPE
Assigned reward Fine-grained process rewards & synthesized DPO preference Dense potential-based rewards from LLM rankings Proxy rewards via latent encoder–decoder
Generalization goal
Interactive env.
Dynamic MARL communication
Table 4: Comparison with other LLM-based credit-assignment methods for non-language agents.
  • Specifically Designed for Language Agents & Interactive Environments: Unlike LCA and LaRe—both devised for classical MARL control domains—our framework is purpose-built for language agents acting in complex UI environments (e.g. mobile navigation, web browsing) that combine large observation spaces with natural-language actions and agent communication.
  • LLM-based process reward assignment: We innovatively employ an LLM as a Critic Agent to assign fine-grained ”process rewards”. The assigned step-level and agent-level process rewards reflect the semantic utility of each action; these rewards are then distilled into synthesized preference data for DPO, moving beyond sparse environmental signals.
  • Emphasis on Cross-Environment Generalization: Our CR strategy is designed to leverage the general world knowledge of LLMs. Combined with continual MARL and mechanisms like communication structure (edge) updates, this significantly enhances the generalization capabilities of language multi-agent systems in unseen environments.
This targeted design delivers superior performance and generalisation for language-driven multi-agent systems, establishing CollabUIAgents as a distinct advancement over contemporaneous credit-assignment methods.
Appendix B Case Studies on Mobile Operation Environments
@@
Appendix C Agentic Fine-Tuning in CollabUIAgents
Figure 3: An example of task execution steps of CollabUIAgentsmobile which is trained on Qwen2 7B and facilitates 4 agents engaging in 3 rounds of conversation. “action_type” represents the action taken, and “index” represents the index of the UI element. The positions of the relevant elements on the UI interface are marked in red.
C.1 Methodology
@@
(1) Basic Environmental Knowledge: This data segment includes identifying UI elements and understanding their properties. We categorize basic knowledge into two types: UI Understanding (coarse-grained): This refers to a broad understanding of the layout and information contained in the UI, such as identifying the purpose of the interface. UI Element Recognition (fine-grained): Since UI typically contains a large number of densely packed interface, the agent needs to be able to distinguish between different types of elements, such as buttons, input fields, and drop-down menus, and understand the associated actions. We develop a series of queries accordingly in Appendix D.4.1, and randomly select UI elements and the layout to assemble queries for the UI Agent.
@@
(3) Process Preference Knowledge: Real-world interactive tasks is quite difficult, and even the most advanced large language model, GPT-4, shows a low task completion rate (30%) in the mobile environment AndroidWorld (Rawles et al., 2025). Training a model solely on scarce successful trajectories still inevitably results in errors. Therefore, as illustrated below Figure 4, we introduce the adversarial agent against the UI agent, and the critic agent to score all actions, obtaining process preference data with step-level rewards. By learning from process preference data, the agent can better distinguish between correct and incorrect actions during the process, ultimately improving task completion rates. The distribution of the collected data can be found in Appendix D.4.2.
@@
C.2 Ablation Study of Agentic Fine-tuning
@@
Appendix D Experiment Details
D.1 Action Space in the Environments
Table 6 and 7 show the action spaces of agents in mobile and web environments, respectively.
@@ D.1 Action Space in the Environments
D.2 Detailed Results in Web Browsing Environments
@@
System #Params/#Agents Input Cross-Task Cross-Website Cross-Domain Avg.
@@
GPT-3.5-Turbo N/A Text 17.4 16.2 18.6 17.4
GPT-4 N/A Text 36.2 30.1 26.4 30.9
@@
Qwen-VL* 9.6B/1A Multimodal 12.6 10.1 08.0 10.2
SeeClick* 9.6B/1A Multimodal 23.7 18.8 20.2 20.9
Qwen2 7B/1A Text 08.6 06.3 07.5 07.4
@@
Base UIAgent 7B/1A Text 13.4 10.6 11.8 11.9
Base UIAgent 7B/4A Text 15.7 11.2 12.9 13.2
CollabUIAgentsmobile 7B/4A Text 19.2 13.8 15.5 16.2
@@
System #Params/#Agents English Chinese Avg.
@@
LLaMA2 7B/1A 03.3 02.5 - - 02.9
LLaMA2 70B/1A 08.3 08.9 - - 10.6
Qwen2 7B/1A 08.6 09.4 08.1 07.8 08.5
@@
Base UIAgent 7B/1A 12.0 13.3 12.7 13.4 12.8
Base UIAgent 7B/4A 13.7 14.5 15.0 13.9 14.0
CollabUIAgentsmobile 7B/4A 18.6 17.7 19.1 15.6 17.7
@@
D.3 Method Implementation
D.3.1 Baselines
@@
(1) M3A (Rawles et al., 2023) is a prompt-based multimodal agent, which combines ReAct- (Yao et al., 2023) and Reflexion-style (Shinn et al., 2023) prompting to interpret user instructions and screen content, then update its decisions. (2) SeeAct (Zheng et al., 2024) is a prompt-based navigation agent originally designed for GPT-4V to perform actions with visual input and textual choices. (3) InfiGUIAgent (Liu et al., 2025a) is used with the temperature set to 0.1, and other settings remain default. (4) DigiRL (Bai et al., 2024) is reproduced on the same dataset as CollabUIAgents, based on Qwen-2.5-VL-3B-Instruct (Wang et al., 2024d). Due to limited computational resource and the high costs of RL training with multiple value estimators, a larger scaled base model could not be applied. (5) SeeClick (Cheng et al., 2024b) is a fine-tuned visual GUI agent that automates tasks relying on screenshots and employs GUI grounding. For those baselines based on closed-source LLMs, we conduct experiments using prompts and settings consistent with Rawles et al. (2025), with model data up to August 2024.
D.3.2 CollabUIAgents Framework
@@
In the agentic fine-tuning process for data collection, we employ GPT-4o-2024-08-06 (OpenAI, 2024) as the UI agent. We set the temperature of different agents to 0.1 to ensure the accuracy of their responses. For the backbone model in the policy multi-agent system, we utilize Qwen2 7B (Yang et al., 2024). Detailed process is depicted in Appendix D.4. In this phase, the model undergoes both supervised learning and off-line preference learning with DPO. Details of the experimental data statistics can be found in Appendix D.4.2. During the supervised fine-tuning phase, the model’s context length is set to 8,192, with a learning rate of 1e-4 and training conducted over 3 epochs. During the preference optimization process, we adapt LoRA fine-tuning (Hu et al., 2022), and the model’s context length is capped at 8,500, with a learning rate of 5e-6. We use bf16 precision, and the training is also carried out for 3 epochs.
@@
In the multi-agent reinforcement learning process, the critic agent is played by GPT-4o-2024-08-06, with its temperature also set to 0.1, providing reward feedback to the rolled out actions. In the MARL phase, the rolling out temperatures of different UI agents are set to 0.1, 0.3, 0.5, and 0.8, respectively. This variation encourages diverse responses and stimulates different behaviors. We use iterative-DPO (Xiong et al., 2024) as the online reinforcement algorithm, and the multi-agent system is updated for each epoch. We train the multi-agent system within the AndroidWorld environment, where the environment seed is randomly generated. Continue MARL on Mind2Web uses instruction datasets autonomously collected with the pipeline in Appendix C. During testing, we maintain consistency with previous work (Rawles et al., 2025) by setting the seed to 30. All our experiments are conducted on 8 A100 80GB GPUs. The model’s context length is capped at 8,500, with a learning rate of 5e-6. We use bf16 precision, and the training is carried out for 3 epochs where the off-policy rolling out is conducted at the beginning of each epoch.
D.4 Data Collection for Agentic Fine-Tuning
@@ 4. What action is associated with UI element X?
D.4.2 Statistics of the Collected Data
@@ D.4.3 Prompts
D.5 Prompts for Different Agents in the CR
@@
D.6 Prompts for CollabUIAgents Framework